# Biomechanics-Aware Trajectory Optimization for Navigation during Robotic Physiotherapy

### v1

@@ Abstract
Robotic devices hold promise for aiding patients in orthopedic rehabilitation. However, current robotic-assisted physiotherapy methods struggle including biomechanical metrics in their control algorithms, crucial for safe and effective therapy. This paper introduces BATON, a Biomechanics-Aware Trajectory Optimization approach to robotic Navigation of human musculoskeletal loads. The method integrates a high-fidelity musculoskeletal model of the human shoulder into real-time control of robot-patient interaction during rotator cuff tendon rehabilitation. We extract skeletal dynamics and tendon loading information from an OpenSim shoulder model to solve an optimal control problem, generating strain-minimizing trajectories. Trajectories were realized on a healthy subject by an impedance-controlled robot while estimating the state of the subject’s shoulder. Target poses were prescribed to design personalized rehabilitation across a wide range of shoulder motion avoiding high-strain areas. BATON was designed with real-time capabilities, enabling continuous trajectory replanning to address unforeseen variations in tendon strain, such as those from changing muscle activation of the subject.
@@
Physical Human-Robot Interaction, Optimization and Optimal Control, Rehabilitation Robotics, Musculoskeletal Dynamics, OpenSim.
@@ I Introduction
Injuries affecting the musculoskeletal system are pervasive in society due to increasing population longevity, strenuous manual labor, and greater engagement in sports [1, 2, 3]. At the forefront of these injuries, disorders impacting the shoulder and the rotator cuff are particularly common, with clinical literature reporting their prevalence to be as high as $22\%$ in the general population [1]. Overall, the demand for therapy and rehabilitation is large and expected to grow, exacerbating the problem posed by a lack of physical therapists. Moreover, the task of treating a complex biological mechanism, such as the shoulder, is challenging due to a fundamental lack of insights into rotator-cuff behavior. Physiotherapists need to move the joint safely, gradually increasing the patient’s range of motion while physically supporting the patient’s arm and avoiding re-injuries [4]. Consequently, simple and limited movements are prescribed, to limit both the risk of re-injury and the physical demands on the therapist to manually manipulate the patient safely.
Robotic technologies can assist in addressing these two gaps during therapy: robots can reduce the manual burden on physiotherapists and provide new tools to monitor and improve rehabilitation outcomes. Overall, rehabilitative robots are already used successfully in post-stroke rehabilitation [5, 6, 7] and gait assistance [8, 9, 10]. However, robots to treat musculoskeletal injuries (such as rotator cuff tears) remain limited. In particular, previous work remains focused on relieving physiotherapists from physically manipulating the patient. Smooth rehabilitation trajectories can be guaranteed by focusing on the control of the movement of the rehabilitative robot itself [11], and therapeutic robotic movements can be learned from expert human demonstrations to automate manual patient manipulation [12]. However, these solutions do not directly account for the patient’s biomechanics. Deeper insights into human musculoskeletal mechanics can enable robots to be aware of tissue and joint loads associated with injury risks, allowing them to operate autonomously or in collaboration with therapists to holistically improve the rehabilitation process [13]. An increasingly promising strategy is to leverage human musculoskeletal models and incorporate them directly into robot control loops [14, 15, 16, 17, 18, 19].
@@
Musculoskeletal modeling has progressed greatly over the past decades, allowing researchers to estimate the activities of individual muscles involved in producing movement [20, 21, 22] and to perform predictive simulations of human motion [23, 24]. As computational models increasingly capture the inner workings of the human body, their applicability continues to expand, particularly in physical human-robot interaction and assistive devices. Recently, biomechanical models have been used to quantify assistance needed by a human operator [14, 15, 16], reduce human metabolic cost in walking-assistive devices [17, 18, 19], and even plan search-and-rescue robotic operations [25] or the motion of supernumerary robotics limbs [26]. Through computer simulations or offline usage of human models, these works demonstrated the importance and utility of including human biomechanics in robot planning and decision-making. However, they did not leverage high-fidelity biomechanical simulations to regulate human-robot interactions in scenarios that require online adaptation to human behavior.
@@
When real-time capabilities are necessary, purely kinematic models have been used to monitor human joint positions, velocities, and torques [27], or to develop controllers for hybrid neuroprostheses of the knee [28]. While lower-fidelity models allow for computational efficiency, they are unsuitable for insights into the musculoskeletal system during shoulder rehabilitation. Biomechanical quantities from a high-fidelity human model were recently used in an online model predictive control framework, achieving predictive control of a leg prosthesis for ergonomically safe walking [29]. Fast predictions of the user’s knee torque were achieved through Bayesian interaction primitives, which could estimate the unobservable biomechanical values via the inclusion of a musculoskeletal model in the data augmentation offline. However, this study did not consider the mechanical behavior of the human tissues (e.g., muscles and tendons).
@@
Building on the previously discussed applications of musculoskeletal models, we can go further to gain insights into the underlying biological tissues that are targets of physical therapy. By estimating tissue loads from external measurements based on a high-fidelity model of the human shoulder, we introduced the concept of “strain maps” [30] to capture the relationship between human pose and the strain (i.e., load) induced in the rotator cuff tendons. Abstracting the rotator cuff tendon strains into an intuitive and navigable map, we achieved offline graph-based motion planning for safe robotic-mediated shoulder rehabilitation [30]. However, the planning algorithm did not account for the dynamics of the human, resulting in trajectories that presented sudden direction changes and were potentially difficult to track. Accelerations imposed on the human shoulder were not monitored, and the trajectories were executed in an open loop. Employing the strain maps, we implemented a reactive impedance-control-based approach to physical human-robot interaction in rehabilitation that enabled a subject to perform therapeutic exercises while robotic assistance protected them from potentially dangerous poses [31]. The system would react only when dangerous movements were already initiated, without anticipating the subject’s trajectory to prevent abrupt movements and experience high corrective forces from the robot.
Other researchers designed optimized trajectories for controlling an ankle rehabilitation robot, to minimize joint loading [32]. A custom ankle model was used to plan the robot’s trajectory offline, making online adjustments based on tracking errors attributed to excessive joint stiffness. However, the human model was not considered during the adjustments, so the reference path could become suboptimal during execution.
To the best of our knowledge, a high-fidelity model of the human musculoskeletal system has never been used to synthesize adaptive robotic rehabilitative movements in real-time through coupled predictive simulations of human dynamics and tissue loading.
Our primary objective is to develop a Biomechanics-Aware Trajectory Optimization for the Navigation of human tissue strains during robotic physiotherapy (BATON). Our novel approach intends to directly couple a state-of-the-art biomechanical model of the human shoulder to robotic control, generating movements for predictive robotic-assisted rotator cuff tendon rehabilitation in real-time. The resulting optimized trajectories are realized by a collaborative robot arm that delivers the movement to a healthy human subject through safe impedance control, with gravity compensation for the human arm naturally and adaptively provided by the model. This allows personalized treatment in the early stages of post-traumatic rehabilitation of the rotator cuff, where the subject does not support the weight of their own arm.
Our specific aims and contributions are:
  • develop an automated system (BATON) to navigate underlying human biomechanical outcomes in real-time during robotic-assisted physiotherapy;
@@
  • design of an optimal control problem (OCP) for trajectory optimization that incorporates a real-time shoulder musculoskeletal model by decoupling skeletal dynamics and strain behavior;
  • analyse performance both in simulation and on a physical robotic system applied to a test subject to demonstrate 1) the effect of OCP parameters on the planned trajectory, 2) the quality of the executed paths, 3) the ability to re-plan online based on changing strain landscapes, 4) real-time computational performance.
An intuitive analogy for navigating around potentially unsafe movements in human rehabilitation is the path planning of an autonomous ship in a bay where underwater reefs are present. To know which maneuvers allow the ship to navigate safely and efficiently toward a target position, we need both a dynamic model of the vessel and a map of the underwater reefs, providing non-obvious insights about the safest path. In our situation, the human shoulder dynamics play the role of the vessel, while the strain maps offer insights into what happens beneath the surface (specifically in the recovering tendons). Only by considering the two elements jointly can safe navigation be obtained.
II Methods
@@
Figure 2: Methodological overview of BATON applied to shoulder physiotherapy. The high-fidelity biomechanical model of the human shoulder (left) is split into skeletal dynamics and rotator cuff tendon loading information (strain maps). An optimal control problem computes safe rehabilitation movements leveraging insights from the human model, and its output is transformed into equivalent references for the robotic controllers to track. The robot’s feedback is used to close the sensing loop on the current human status.
Our approach joins high-fidelity human biomechanics, trajectory optimization, and robot control to generate safe rehabilitation trajectories based on real-time underlying biomechanical insights (Fig. 2). We present our method for extracting insights from the musculoskeletal model for real-time use in optimization and control in Section II-A. Low-strain rehabilitation movements are planned towards target human poses by solving a trajectory optimization problem, detailed in Section II-B. A collaborative KUKA LBR iiwa 7 robotic arm is employed to estimate the current human pose (Section II-D) and to deliver the actual rehabilitation movement to the human subject during therapy. Details about the robot’s control are given in Section II-E, and our experimental design is presented in Section II-F.
@@ II-A Human Biomechanical Model
We employ a high-fidelity biomechanical model of the human shoulder [33] to capture the human response in a computer simulation and predict how movements and forces will affect the tendons’ strain in the rotator cuff. The human model, developed in OpenSim [34, 35], captures the shoulder’s movement degrees of freedom (DoFs) and the muscles powering it. As such, it can inform the robot’s controller of the internal tissue and skeletal mechanics during the physical human-robot interaction. In particular, we are interested in measures related to injury risk of the rotator cuff muscles (i.e., supraspinatus, subscapularis, infraspinatus, and teres minor), which act as stabilizers for the glenohumeral joint in the shoulder and whose tendons are subject to tears due to overuse. Mechanical strain is associated with re-injury risk of rotator cuff tendons, as excessive stretching of a healing tendon could lead to damages [36].
@@
In the human body, the rotator cuff spans the glenohumeral joint, which permits mobility of the upper arm (humerus) with respect to the shoulder blade (scapula) through 3 DoFs. As such, we extracted a reduced-order model of the human shoulder, capturing the mobility of this joint alone (see Fig. 2, left). The state of the resulting human model is described as:
$\boldsymbol{x}=[PE,\dot{PE},SE,\dot{SE},AR,\dot{AR}]^{T}$ (1)
where $PE,SE$ and $AR$ are the 3 movement DoFs defined as the Y-X’-Y” sequence of intrinsic Euler angles in the fixed shoulder reference frame ($PE$: plane of elevation, $SE$: shoulder elevation as rotation around -X, $AR$: axial rotation, visualized in Fig. 3), and $\dot{PE},\dot{SE}$ and $\dot{AR}$ are their corresponding derivatives with respect to time. The arm size and mass of our human subject were used to personalize the model’s properties through OpenSim’s scaling functionalities and tabulated anthropometric data [37].
@@
The personalized shoulder model specifies how the shoulder moves under the influence of applied torques/forces, and how these affect the underlying musculoskeletal system. However, embedding the model directly to plan rehabilitation movements through the OpenSim API is not computationally tractable for real-time applications. To achieve real-time performances without sacrificing model accuracy, we decouple tissue mechanics describing tendon strain changes with shoulder movement from skeletal dynamics that simulate how shoulder motion is generated as a consequence of applied forces. This decoupling enables us to retain the information contained in the original OpenSim model and access it much more rapidly, to formulate the problem of navigating shoulder motion over strain maps governed by skeletal dynamics.
II-A1 Generating differentiable strain maps
the musculoskeletal model can be queried to provide quantitative information on the mechanical strain $\sigma$ of a desired tendon in the rotator cuff. Biologically, the model’s state $\boldsymbol{x}$ and the activation level $\alpha\in[0,1]$ of the corresponding muscle determine tendon strain, which can be expressed as:
$\sigma=\frac{l_{T}(\boldsymbol{x},\alpha)-l_{0}}{l_{0}}\cdot 100\%.$ (2)

### v2

# Biomechanics-Aware Trajectory Optimization for Online Navigation during Robotic Physiotherapy

@@ Abstract
Robotic devices provide a great opportunity to assist in delivering physical therapy and rehabilitation movements, yet current robot-assisted methods struggle to incorporate biomechanical metrics essential for safe and effective therapy. We introduce BATON, a Biomechanics-Aware Trajectory Optimization approach to online robotic Navigation of human musculoskeletal loads for rotator cuff rehabilitation. BATON embeds a high-fidelity OpenSim model of the human shoulder into an optimal control framework, generating strain-minimizing trajectories for real-time control of therapeutic movements. Its core strength lies in the ability to adapt biomechanics-informed trajectories online to unpredictable volitional human actions or reflexive reactions during physical human-robot interaction based on robot-sensed motion and forces. BATON’s adaptability is enabled by a real-time, model-based estimator that infers changes in muscle activity via a rapid redundancy solver driven by robot pose and force/torque sensor data. We validated BATON through physical human-robot interaction experiments, assessing response speed, motion smoothness, and interaction forces.
@@
Physical Human-Robot Interaction, Optimization and Optimal Control, Rehabilitation Robotics, Musculoskeletal Dynamics.
@@ I Introduction
Injuries affecting the musculoskeletal system are pervasive in society due to an increasingly aging population, strenuous manual labor, and more widespread engagement in sports [1, 2, 3]. Disorders impacting the shoulder and the rotator cuff are particularly common, with clinical literature reporting their prevalence to be as high as $22\%$ in the general population [1] and 50% over age 66 [4]. Overall, the demand for therapy and rehabilitation is large and expected to grow, exacerbating the problem posed by a lack of physical therapists [5]. Moreover, the task of treating a complex biological mechanism, such as the shoulder, is challenging due to a fundamental lack of insights into rotator cuff behavior during rehabilitation. Physiotherapists need to move the joint safely, gradually increasing the patient’s range of motion while physically supporting the patient’s arm and avoiding re-injuries [6]. Consequently, simple and limited movements are prescribed, to limit both the risk of re-injury and the physical demands on the therapist to manually manipulate the patient safely. However, overly conservative movements may result in a limited range of motion during the therapy, which can delay recovery since moving through a larger range of motion was shown to expedite recovery [7, 8]
Robotic technologies can assist in addressing manual manipulation and re-injury risks during therapy: robots can reduce the manual burden on physiotherapists and provide new tools to monitor and improve rehabilitation outcomes. Overall, rehabilitative robots are already used successfully in post-stroke rehabilitation [9, 10, 11] and gait assistance [12, 13, 14]. However, robots to treat musculoskeletal injuries (such as rotator cuff tears) remain limited. In particular, previous work remains focused on relieving physiotherapists from physically manipulating the patient. Smooth rehabilitation trajectories can be guaranteed by focusing on the control of the movement of the rehabilitative robot itself [15], and therapeutic robotic movements can be learned from expert human demonstrations to automate manual patient manipulation [16]. However, these solutions do not directly account for the patient’s biomechanics. Deeper insights into human musculoskeletal mechanics can enable robots to be aware of tissue and joint loads associated with injury risks, allowing them to operate autonomously or in collaboration with therapists to holistically improve the rehabilitation process [17]. An increasingly promising strategy is to leverage human musculoskeletal models and incorporate them directly into robot control loops [18, 19, 20, 21, 22, 23, 24].
@@
Musculoskeletal modeling has progressed greatly over the past decades, allowing researchers to estimate the activities of individual muscles involved in producing movement [25, 26, 27] and to perform predictive simulations of human motion [28, 29]. As computational models increasingly capture the inner workings of the human body, their applicability continues to expand, particularly in physical human-robot interaction and assistive devices. Recently, biomechanical models have been used to quantify assistance needed by a human operator [18, 19, 20], reduce human metabolic cost in walking-assistive devices [21, 22, 23], and even plan search-and-rescue robotic operations [30] or the motion of supernumerary robotics limbs [31]. Other researchers designed optimized trajectories for controlling an ankle rehabilitation robot, to minimize joint loading [32]. A custom ankle model was used to plan the robot’s trajectory offline, making online model-free adjustments based on tracking errors attributed to excessive joint stiffness. Through computer simulations or offline usage of human models, these works demonstrated the importance and utility of including human biomechanics in robot planning and decision-making. However, they did not leverage high-fidelity biomechanical simulations to regulate human-robot interactions in scenarios that require online adaptation to human behavior, such as robotic physiotherapy.
@@
When real-time capabilities are necessary, purely kinematic models have been used to monitor human joint positions, velocities, and torques [33], or to develop controllers for hybrid neuroprostheses [34]. While lower-fidelity models allow for computational efficiency, they are unsuitable for deep insights into the musculoskeletal system during physical human-robot interaction.
Recently, biomechanical quantities from a high-fidelity human model have been used in an online model predictive control framework, achieving predictive control of a leg prosthesis for ergonomically safe walking [35]. Researchers leveraged fast predictions of how knee torque relates to ankle angle, by including offline information from a musculoskeletal model in an imitation learning strategy. Moya-Esteban and Sartori, instead, directly leveraged surface EMG-driven human models to minimize lumbar joint compression during external load handling, modulating the assistance of a wearable back exoskeleton online [36, 37].
However, these studies did not consider the mechanical behavior of the human tissues (e.g., muscles and tendons) during physical human-robot interaction for shoulder injuries. Without such consideration, safety cannot be ensured throughout the rehabilitation movements. Furthermore, surface EMG is limited to the measurement of surface muscles, while the underlying muscles and tissues cannot be monitored. Since typical shoulder injuries involve the rotator cuff muscles, which are deep-lying muscles that hold the arm in place, there is a need for insights into the underlying biological tissues.
@@
To gain such underlying insights, tissue loads can be estimated from external measurements based on a high-fidelity model of the human shoulder. In this direction, researchers developed the concept of “strain maps” [38] to intuitively represent musculoskeletal tissue loadings. A strain map captures the relationship between human pose and the strain (i.e., load) induced in the rotator cuff tendons. By abstracting the rotator cuff tendon strains into an intuitive and navigable map, offline graph-based motion planning for safe robotic-mediated shoulder rehabilitation was achieved. However, the planning algorithm did not account for the dynamics of the human, resulting in trajectories that presented sudden direction changes, potentially difficult to track, and dynamically unsafe.
Moreover, the method in [38] did not consider online variations in tendon strain level caused by muscle activations during therapy, limiting its real-world applicability. Employing the strain maps, a reactive impedance-control-based approach to physical human-robot interaction in rehabilitation was also presented, to enable a subject to perform therapeutic exercises while robotic assistance protected them from navigating through potentially dangerous poses [39]. In this case, the system would react only when dangerous movements were already initiated, without anticipating the subject’s trajectory, which could lead to abrupt movements and high corrective forces from the robot. Critically, unpredictable volitional actions of the patient and/or reflexive reactions to physical human-robot interaction in terms of muscle activity and their effects on tendon strains were not considered.
It is clear that musculoskeletal models have the potential to inform both the planning of the therapeutic motion, in regard to tissue loading, and the human activity and response to robot movement through physical human-robot interaction. However, there has been no single framework for combining planning, execution, and human response during robotic rehabilitation in an online manner.
To address this gap, our overarching objective is to develop a Biomechanics-Aware Trajectory Optimization for the Navigation of real-time human tissue strains during robotic physiotherapy (BATON) targeting musculoskeletal rehabilitation. Our novel approach directly couples a state-of-the-art biomechanical model of the human shoulder to robotic control, generating movements for predictive robotic-assisted musculoskeletal rehabilitation in real-time. This is achieved by accounting for changes in muscle activity due to unpredictable volitional actions of the patient and/or reflexive reaction to physical human-robot interaction that is estimated directly from robot sensors. The resulting planned trajectories are realized by a collaborative robot arm that delivers the movement to a healthy human subject through safe impedance control and provides estimations of contact forces and human poses to close the loop. We propose a muscle activity estimation that provides an alternative to sensor-based muscle activity estimation from surface EMG. Our system can inform the online path planning about the activity of deep-lying muscles, such as rotator cuffs, which cannot be measured with surface EMG.
To achieve our objective, we target the following specific aims:
@@
  • A1: design a biomechanics-aware trajectory optimization planner that efficiently incorporates a real-time shoulder musculoskeletal model into the underlying optimal control problem (OCP) by decoupling skeletal dynamics and tendon strain behavior of the rotator cuff muscles;
  • A2: develop a method to inform the planner of the dynamic changes in human biomechanics dependent on muscle activity during real-time physical human-robot interaction;
  • A3: experimentally evaluate the performance of BATON in terms of the speed, smoothness, and forces of human-in-the-loop response.
@@
Figure 2: Methodological overview of BATON applied to physiotherapy for the rotator cuff muscle tendons of the human shoulder. The high-fidelity biomechanical model of the human shoulder (left), employed fully to estimate human muscle activations, is decoupled in skeletal dynamics and rotator cuff tendon loading information (strain maps) for efficient trajectory planning. An optimal control problem computes safe rehabilitation trajectories leveraging insights from the human model, and its output is transformed into equivalent references for the robotic controllers to track during physical human-robot interaction (pHRI). The robot’s feedback on the current human status is used to close the loop, allowing real-time estimation of the physiological activation for the muscles of interest.
Our method, BATON (see Fig. 2, for overview) extracts insights from the musculoskeletal model for real-time use in optimization and control (Sec. II-A). To exercise the range of motion of the human shoulder during therapy, the planner finds the minimum strain path by solving a trajectory optimization problem (Sec. II-B). However, the physical interaction between the robot and the human during trajectory execution induces loads that the human responds to voluntarily and/or reflexively. Since human decisions, sensitivities, and responses to movement and other stimuli are inherently unpredictable, it is imperative that the human response in the control loop is based on actual measurements. Such human-in-the-loop actions and reactions change the level of muscle activity. Thus, muscle forces and corresponding tissue loads cannot be fully predicted, but need to be incorporated. To this end, we estimate these actions and reactions online using a state-of-the-art muscle redundancy solver (Sec. III-A). We employ a collaborative robotic arm instrumented with a force-torque sensor to estimate the current human pose and loading (Sec. III-B) and administer rehabilitation movements to the human subject. We embody the robot controller in a closed-loop with the subject (Sec. IV-A) to perform physical experiments (Sec. IV-B) that evaluate our improvements over the state-of-the-art rotator cuff rehabilitation techniques.
II Motion Planning on Strain Maps
Musculoskeletal insights are derived from a human biomechanical model, becoming the strain maps navigated by our online trajectory planner. We developed a rehabilitation path planner to traverse these maps to synthesize safe robotic-led human movements by formulating and efficiently solving an optimal control problem.
To give a practical analogy of how we navigate safety during rehabilitation, let us consider the path planning of a ship in a bay where underwater reefs are present. To know which maneuvers allow the ship to navigate safely and efficiently toward a target position, we need both a dynamic model of the vessel and a map of the underwater reefs, providing non-obvious insights about the safest path. In our situation, the human shoulder dynamics play the role of the vessel, while the strain maps offer insights into what happens beneath the surface (in the tissue being rehabilitated). Safe navigation can only be obtained by considering where it is safe to travel, together with the maneuverability of the ship.
@@ II-A Human Biomechanical Model
We employ a high-fidelity biomechanical model of the human shoulder [40], developed in OpenSim [41, 42], to capture unpredictable volitional actions of the human and/or responses to motions and forces induced by robot interactions. The human model captures the shoulder’s movement degrees of freedom (DoFs) and the muscles that generate movement. The model, therefore, can inform the robot’s controller of the internal human mechanics during the physical human-robot interaction. In particular, we are interested in quantities related to injury risk of the rotator cuff muscles (i.e., supraspinatus, subscapularis, infraspinatus, and teres minor), which act as stabilizers for the glenohumeral (GH) joint in the shoulder and whose tendons are subject to tears due to overload, chronic weakness, and impingement. With the model, we predict how movements and forces affect tendon strain in the rotator cuffs, which cannot be measured by surface EMG due to their position deep inside the body. Mechanical strain is associated with re-injury risk of rotator cuff tendons, as excessive stretching of a healing tendon could lead to tearing [43].
@@
In the human body, the rotator cuff spans the GH joint, which permits mobility of the upper arm (humerus) with respect to the shoulder blade (scapula) through 3 DoFs. To improve efficiency without sacrificing fidelity, we reduced the original model to the 3 DoFs and 22 muscles that actuate the GH joint only, to isolate the effects on the rotator cuff muscles (see Fig. 2, left). The arm dimensions and mass of our human subject were used to personalize the model’s properties through OpenSim’s scaling functionalities and tabulated anthropometric data [44]. We define the state of the resulting human model as:
$\boldsymbol{\theta}=[PE,SE,AR]^{T}$ (1)
$\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},...,\alpha_{22}]^{T},\ \alpha_{i}\in[0,1]\ \forall i$ (2)
$\boldsymbol{x}=[\boldsymbol{\theta}^{T},\dot{\boldsymbol{\theta}}^{T},\boldsymbol{\alpha}^{T}]^{T}$ (3)
where $PE,SE$, and $AR$ are the 3 movement DoFs defined as the Y-X’-Y” sequence of intrinsic Euler angles in the fixed shoulder reference frame ($PE$: plane of elevation, $SE$: shoulder elevation as rotation around -X, $AR$: axial rotation, visualized in Fig. 3 and indicated collectively as $\boldsymbol{\theta}\in\mathbb{R}^{3}$), $\dot{\boldsymbol{\theta}}\in\mathbb{R}^{3}$ is the vector containing their derivatives with respect to time, and $\boldsymbol{\alpha}$ contains the muscle activation level for each of the muscles in the model. Note that, in the remainder of this paper, we will refer to $\boldsymbol{\mathrm{q}}=[\boldsymbol{\theta}^{T},\dot{\boldsymbol{\theta}}^{T}]^{T}$ as the purely kinematic state of the model, while $\boldsymbol{x}$ is the complete musculoskeletal model state.
@@
The personalized shoulder model specifies how the shoulder moves under the influence of externally applied torques/forces, and how these affect the underlying musculoskeletal system. However, embedding the model directly to plan rehabilitation movements through the OpenSim API is not computationally tractable for real-time applications. To achieve real-time performances without sacrificing model accuracy, we decouple tissue mechanics that describe tendon strain with shoulder movement and muscle activation from skeletal dynamics that simulate how passive shoulder motion is generated as a consequence of externally applied forces. This decoupling enables us to formulate the problem of navigating shoulder motion over strain maps governed by skeletal dynamics and to solve planning efficiently.
II-A1 Navigable strain maps
A strain map describes how strains in individual tendons, or collectively (e.g. the peak), change with respect to the model state (motion and activation). As the complete human model state $\boldsymbol{x}$ spans a high-dimensional space, we focus on the pose (target position from the therapist) and overall level of activation as the primary dimensions of a spatial strain map. We compute the strain map querying the musculoskeletal model to extract quantitative information on the mechanical strain $\sigma_{i}$ of every ith tendon in the rotator cuff. Biologically, tendon strain depends on the model’s state $\boldsymbol{x}$ as:
$\sigma_{i}=\frac{l_{i}(\boldsymbol{x})-l_{i,0}}{l_{i,0}}\cdot 100\%.$ (4)